# thanks: Present address: SLAC National Laboratory

### v4

@@ II Theoretical model
$-16\left(\left(\frac{\Delta\omega}{\omega_{s}}\right)^{2}+\left(\frac{\alpha}{\omega_{s}}\right)^{2}\right)\psi^{2}_{\tau}=0.$
@@ III Experiment
To carry out the RF phase modulation at KARA, we have introduced the modulation function into our low-level RF (LLRF) system (Dimtel, LLRF9/500 [7]). Each RF station in the KARA storage ring has one klystron, whose maximum output power is 250 kW. Because one RF station has two RF accelerating cavities, one klystron drives two RF cavities in KARA. Each RF station has its own LLRF system to control two RF stations independently. From one master oscillator, whose signal frequency is around 499.7 MHz, each LLRF system in each RF station receives the master clock signal and drives the RF station. Because the LLRF system has the function of phase modulation, we can perform phase modulation for each RF station independently. In the experiments discussed in the following part, we performed phase modulation at only one RF station, which is the same condition as in the simulation discussed in section IV. To perform the phase modulation, we remotely set and measure the modulation frequency and amplitude using the LLRF system. The LLRF system has a function for measuring the modulation pattern of the forward power to the cavities. By analyzing the modulation pattern, we measured the modulation amplitude and frequency. The phase modulation amplitude values we applied in the experiments were $10.25\pm 0.37$, $13.45\pm 0.56$, and $16.42\pm 1.36$ degrees, with the errors in the modulation phase arising from reproducibility. As discussed in section IV, the simulation results agree with the experiments and show slightly smaller modulation phase amplitudes than the values obtained from the forward power measurement. This suggests the possibility that the modulation amplitude in the RF cavity is smaller than that in the forward power. The technical details for the method of phase modulation by the LLRF system are discussed in [8].
@@ III.1 Modulation frequency and amplitude
To discuss the negative detuning character with the measured detuning curve in depth, we measured the synchrotron oscillation frequency $f_{s}$ in parallel with the bunch length measurement by measuring the longitudinal frequency spectrum using the bunch-by-bunch feedback system [11]. The result of the measured synchrotron frequency was $(32.6\pm 0.54)$ kHz, where the measurement error is due to the uncertainties in the frequency readout. This meant that the detuning curves in Fig. 4, even though in the most prominent negative detuning case with 16.42 degrees of modulation, did not show the negative detuning character discussed in section II. However, we found that the readout value of the bunch-by-bunch feedback system had a systematic error by comparing the measured synchrotron frequency with a phase tracking method [12, 13], which allows us to measure the synchrotron frequency precisely. Therefore, we conducted another measurement to evaluate the systematic difference in the synchrotron frequency between with and without phase tracking. Figure 5 shows a typical longitudinal spectrum measured with the bunch-by-bunch feedback system with and without phase tracking. The beam spectrum has a broad peak, which we analyzed as the synchrotron oscillation spectrum and is indicated by the blue ellipse in the figure, and a sharp peak indicated by the arrow in the figure that always appears in a higher frequency region than the broad peak when phase tracking is on. We analyzed the difference in frequency between the broad and sharp peaks and treated it as a systematic error in the measurement of the synchrotron frequency. We repeated the measurement with several different RF accelerating voltages and beam current conditions while keeping a multi-bunch and 2.5 GeV condition. From the measurement, the systematic error of the measured synchrotron frequency was evaluated to be $(+1.73\pm 0.36)$ kHz. From the estimation, the double synchrotron frequency, including the systematic error, was evaluated as $(68.6\pm 1.8)$ kHz. The detuning curve of 16.42 degrees in Fig. 4 peaks around 66.6 kHz, indicating negative detuning from the double synchrotron oscillation period, considering the error. As discussed in section II, the negative detuning frequencies, assuming the phase-modulated bunch length lengthens to $\sqrt{2}$ times the natural bunch length and the maximum bunch length with a modulation amplitude of 16.42 degrees in Fig. 4, are estimated to be around -0.4 and -0.9 kHz, which are smaller than the readout error value of the synchrotron frequency in this experiment. The largest modulation phase amplitude case in this experiment indicates only the observable negative frequency detuning.
@@ IV.1 Modulation frequency and amplitude
Figure 8 shows the simulation results of longitudinal phase space snapshots in the case of active RF phase modulation, where the modulation amplitude is 7.91 degrees and the modulation frequency is adjusted to maximize the bunch length condition of $12002f_{s}-$ 1200 Hz. The total beam current is set at 90 mA in the simulation. The time difference between these two snapshots in Fig. 8 is a quarter of one synchrotron oscillation period. These results show that both the bunch length and the energy spread change with phase modulation. Two spiral arms emerge from the beam core, indicating nonlinear effects.
@@
Figure 8: The simulation result of the longitudinal phase space distribution when the RF phase modulation is active with a total beam current of 90 mA. The modulation frequency and amplitude are set to $12002f_{s}-$ 1200 Hz and 7.91 degrees, respectively. The blue and red dots correspond to snapshots of the electron distribution with a time difference of a quarter of one synchrotron oscillation period.
Figure 9: The simulation results of the change in the longitudinal electron distribution in one bunch when the modulation frequency is set to maximize the bunch lengthening [$12002f_{s}-$ 1200 Hz;(a)], -1 kHz, and +1 kHz detuning from condition (a) [(b), (c)]. The 90 mA total beam current is set in the simulation, and all the other machine parameters are the same as in Fig. 8. The vertical scale measures the bunch length. The positive direction on the vertical axis corresponds to the head of the bunch. The blue and yellow colors indicate low and high electron density in the charge distribution, respectively.

### v5

# Detuning Properties of RF Phase Modulation in an Electron Storage Ring

@@ II Theoretical model
[…]
@@ III Experiment
To carry out the RF phase modulation at KARA, we have introduced the modulation function into our low-level RF (LLRF) system (Dimtel, LLRF9/500 [7]). Each RF station in the KARA storage ring has one klystron, whose maximum output power is 250 kW. Because one RF station has two RF accelerating cavities, one klystron drives two RF cavities in KARA. Each RF station has its own LLRF system to control two RF stations independently. From one master oscillator, whose signal frequency is around 499.7 MHz, each LLRF system in each RF station receives the master clock signal and drives the RF station. Because the LLRF system has the function of phase modulation, we can perform phase modulation for each RF station independently. In the experiments discussed in the following part, we performed phase modulation at only one RF station, which is the same condition as in the simulation discussed in section IV. To perform the phase modulation, we remotely set and measure the modulation frequency and amplitude using the LLRF system. The LLRF system has a function for measuring the modulation pattern of the forward power to the cavities. By analyzing the modulation pattern, we measured the modulation amplitude and frequency. The phase modulation amplitude values we applied in the experiments were 10.25±0.3710.25$\pm$ 0.3710.25 ± 0.37, 13.45±0.5613.45$\pm$ 0.5613.45 ± 0.56, and 16.42±1.3616.42$\pm$ 1.3616.42 ± 1.36 degrees, with the errors in the modulation phase arising from reproducibility. As discussed in section IV, the simulation results agree with the experiments and show slightly smaller modulation phase amplitudes than the values obtained from the forward power measurement. This suggests the possibility that the modulation amplitude in the RF cavity is smaller than that in the forward power. The technical details for the method of phase modulation by the LLRF system are discussed in [8].
@@ III.1 Modulation frequency and amplitude
To discuss the negative detuning character with the measured detuning curve in depth, we measured the synchrotron oscillation frequency $f_{s}$ in parallel with the bunch length measurement by measuring the longitudinal frequency spectrum using the bunch-by-bunch feedback system [11]. The result of the measured synchrotron frequency was (32.6±0.54)(32.6$\pm$ 0.54)( 32.6 ± 0.54 ) kHz, where the measurement error is due to the uncertainties in the frequency readout. This meant that the detuning curves in Fig. 4, even though in the most prominent negative detuning case with 16.42 degrees of modulation, did not show the negative detuning character discussed in section II. However, we found that the readout value of the bunch-by-bunch feedback system had a systematic error by comparing the measured synchrotron frequency with a phase tracking method [12, 13], which allows us to measure the synchrotron frequency precisely. Therefore, we conducted another measurement to evaluate the systematic difference in the synchrotron frequency between with and without phase tracking. Figure 5 shows a typical longitudinal spectrum measured with the bunch-by-bunch feedback system with and without phase tracking. The beam spectrum has a broad peak, which we analyzed as the synchrotron oscillation spectrum and is indicated by the blue ellipse in the figure, and a sharp peak indicated by the arrow in the figure that always appears in a higher frequency region than the broad peak when phase tracking is on. We analyzed the difference in frequency between the broad and sharp peaks and treated it as a systematic error in the measurement of the synchrotron frequency. We repeated the measurement with several different RF accelerating voltages and beam current conditions while keeping a multi-bunch and 2.5 GeV condition. From the measurement, the systematic error of the measured synchrotron frequency was evaluated to be (+1.73±0.36)(+1.73$\pm$ 0.36)( + 1.73 ± 0.36 ) kHz. From the estimation, the double synchrotron frequency, including the systematic error, was evaluated as (68.6±1.8)(68.6$\pm$ 1.8)( 68.6 ± 1.8 ) kHz. The detuning curve of 16.42 degrees in Fig. 4 peaks around 66.6 kHz, indicating negative detuning from the double synchrotron oscillation period, considering the error. As discussed in section II, the negative detuning frequencies, assuming the phase-modulated bunch length lengthens to $\sqrt{2}$ times the natural bunch length and the maximum bunch length with a modulation amplitude of 16.42 degrees in Fig. 4, are estimated to be around -0.4 and -0.9 kHz, which are smaller than the readout error value of the synchrotron frequency in this experiment. The largest modulation phase amplitude case in this experiment indicates only the observable negative frequency detuning.
@@ IV.1 Modulation frequency and amplitude
Figure 8 shows the simulation results of longitudinal phase space snapshots in the case of active RF phase modulation, where the modulation amplitude is 7.91 degrees and the modulation frequency is adjusted to maximize the bunch length condition of $2f_{s}-1200$ Hz. The total beam current is set at 90 mA in the simulation. The time difference between these two snapshots in Fig. 8 is a quarter of one synchrotron oscillation period. These results show that both the bunch length and the energy spread change with phase modulation. Two spiral arms emerge from the beam core, indicating nonlinear effects.
@@
Figure 8: The simulation result of the longitudinal phase space distribution when the RF phase modulation is active with a total beam current of 90 mA. The modulation frequency and amplitude are set to $2f_{s}-1200$ Hz and 7.91 degrees, respectively. The blue and red dots correspond to snapshots of the electron distribution with a time difference of a quarter of one synchrotron oscillation period.
Figure 9: The simulation results of the change in the longitudinal electron distribution in one bunch when the modulation frequency is set to maximize the bunch lengthening [$2f_{s}-1200$ Hz;(a)], -1 kHz, and +1 kHz detuning from condition (a) [(b), (c)]. The 90 mA total beam current is set in the simulation, and all the other machine parameters are the same as in Fig. 8. The vertical scale measures the bunch length. The positive direction on the vertical axis corresponds to the head of the bunch. The blue and yellow colors indicate low and high electron density in the charge distribution, respectively.